OpenStudy (anonymous):

Oil spilled from a ruptured tanker spreads in a circle whose area increases at a constant rate of 9 mi^2/hr . How rapidly is radius of the spill increasing when the area is 10 mi^2 ?

OpenStudy (anonymous):

3 mi/hr (square root of 9, because square of radius increases in proportion to area)

OpenStudy (anonymous):

wait, I'm on crack.

OpenStudy (anonymous):

me too.

OpenStudy (anonymous):

me too.

OpenStudy (anonymous):

\[A=\pi r^2\] \[A'=2 \pi r r'\] \[r'=\frac{A'}{2\pi r}\]

OpenStudy (anonymous):

\[A'=9\] so \[r'=\frac{9}{2\pi r}\]

OpenStudy (anonymous):

if \[A=10\] then \[r=\sqrt{\frac{10}{\pi}}\]

OpenStudy (anonymous):

plug that in and get the answer \[r'=\frac{9}{2\pi}\times \sqrt{\frac{\pi}{10}}\]

OpenStudy (anonymous):

im lost is that the answer you got? because when i pluged it in it said it was wrong

OpenStudy (anonymous):

A = pi*r^2 dA/dt = 2*pi*r*dr/dt dr/dt = A'/(2*pi*r) = 9/(2*pi*10) = 9/20pi = 0.1432

OpenStudy (anonymous):

oh wait...no

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

satellite is right

OpenStudy (anonymous):

you are told area is 10 yes? and \[A=\pi r^2\] making \[r=\sqrt{\frac{10}{\pi}}\]

OpenStudy (anonymous):

you obviously entered it into the calculator incorrectly

OpenStudy (anonymous):

do you want a decimal for this?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

= 0.8029

OpenStudy (anonymous):

that is what i get as well

OpenStudy (anonymous):

yea that was perfect thanks